# Co-Seg++: Mutual Prompt-Guided Collaborative Learning for Versatile Medical Segmentation

### v1

@@ Abstract
Medical image analysis is critical yet challenged by the need of jointly segmenting organs or tissues, and numerous instances for anatomical structures and tumor microenvironment analysis. Existing studies typically formulated different segmentation tasks in isolation, which overlooks the fundamental interdependencies between these tasks, leading to suboptimal segmentation performance and insufficient medical image understanding. To address this issue, we propose a Co-Seg++ framework for versatile medical segmentation. Specifically, we introduce a novel co-segmentation paradigm, allowing semantic and instance segmentation tasks to mutually enhance each other. We first devise a spatio-temporal prompt encoder (STP-Encoder) to capture long-range spatial and temporal relationships between segmentation regions and image embeddings as prior spatial constraints. Moreover, we devise a multi-task collaborative decoder (MTC-Decoder) that leverages cross-guidance to strengthen the contextual consistency of both tasks, jointly computing semantic and instance segmentation masks. Extensive experiments on diverse CT and histopathology datasets demonstrate that the proposed Co-Seg++ outperforms state-of-the-arts in the semantic, instance, and panoptic segmentation of dental anatomical structures, histopathology tissues, and nuclei instances. The source code is available at https://github.com/xq141839/Co-Seg-Plus.
@@ 1 Introduction
Medical image segmentation plays a crucial role in clinical applications and has received significant research attention. [1, 2, 3]. For example, histopathology image analysis demands accurate segmentation of tissue regions and separation of individual nuclei within these regions to assess tissue subtypes and tumor grading [4]. Similarly, cone beam computed tomography (CBCT) requires precise delineation of anatomical structures such as jaw bones and sinuses, alongside instance-level segmentation of individual teeth for dental treatment planning [5]. These dual requirements have given rise to the challenging field of versatile medical segmentation, which seeks to unify both semantic and instance segmentation tasks across various medical imaging contexts.
@@
Typically, medical image segmentation frameworks [6, 7, 8] have been designed for specific tasks, resulting in isolated solutions that address semantic segmentation and instance segmentation separately, as shown in Fig. 1(a). These conventional approaches typically adopt independent encoder-decoder architectures without shared information during the training process. The semantic segmentation networks are optimized to classify each pixel into pre-defined categories, meanwhile instance segmentation models [9, 10, 2] focus on detecting and delineating individual objects. However, multiple independent networks [11, 12] lead to increased computation complexity. More seriously, this separation learning [13] impairs consistency for similar low-level features across tasks, resulting in redundant feature extraction. These limitations motivate the development of more integrated approaches that could leverage commonalities between semantic and instance segmentation tasks in medical images.
@@
To overcome this bottleneck, we propose Co-Seg++, a collaborative versatile medical segmentation framework that allows semantic and instance segmentation tasks to mutually enhance each other. As illustrated in Fig. 1(c), Co-Seg++, based on a novel co-segmentation paradigm, improves segmentation mask quality by capturing contextual dependencies between the two tasks. Specifically, we design a spatio-temporal prompt encoder (STP-Encoder) that adaptively integrates spatial context and temporal dynamics to generate prior spatial constraints for guiding segmentation decoding across both tasks. We further introduce a multi-task collaborative decoder (MTC-Decoder) that leverages cross-task guidance to jointly predict semantic and instance segmentation masks in a unified framework. This collaborative learning strategy not only achieves enhanced contextual consistency but also significantly reduces prediction errors through mutual task optimization. Extensive experiments on diverse histopathology and dental datasets demonstrate that our Co-Seg++ consistently outperforms state-of-the-art segmentation methods.
@@
  • We devise an STP-Encoder that integrates long-range spatial and temporal information to extract high-quality semantic and instance prompts, providing effective prior knowledge for cross-task segmentation guidance.
@@
Figure 3: The overview of our Co-Seg++ framework for collaborative versatile medical segmentation. For ease of understanding, we elaborate on the case of Co-Seg++ in tissue and nuclei segmentation. Our Co-Seg++ fully exploits complementary context by leveraging spatio-temporal prompts.
@@ 2.1 Medical Image Segmentation
Medical image segmentation is fundamental to computer-aided diagnosis and treatment planning [6, 7, 8]. This field encompasses semantic segmentation (i.e., pixel-level classification), instance segmentation (i.e., individual object delineation), and panoptic segmentation that unifies both by jointly predicting semantic classes and instance identities [18]. Building on the U-Net encoder-decoder framework [6], recent advances span three network architectures: convolutional neural network (CNN), vision transformers (ViT) [19], and more recent Mamba models [20]. The CNN-based methods [7, 21, 8, 22] applied automated architecture optimization and multi-scale convolutions to enhance target awareness. The ViT-based methods [23, 24] leveraged self-attention mechanisms to capture global contexts. The Mamba-based methods [12, 25, 26] utilized sequence modeling to handle long-range relationships with linear computational complexity. These works have demonstrated strong adaptability across diverse imaging modalities.
@@ 2.2 Versatile Segmentation Paradigms
The classical versatile medical segmentation frameworks have predominantly adopted a task-decoupled paradigm to address multiple segmentation objectives simultaneously. The TIAToolbox [13] employed an ensemble approach that integrates multiple independent networks for tissue and nuclei segmentation in histopathology images, treating each task as a separate optimization problem. To enhance cross-task consistency, Cerberus [14] employed the shared encoder with decoupled decoders for joint training efficiency. Federico et al. [31] adopted multi-class strategies with independent labels for CBCT tasks. However, they lack flexible guidance mechanisms that limit their ability to incorporate contextual knowledge or interactive signals, highlighting the need for more adaptive and guidance-driven segmentation paradigms.
The segment anything model (SAM) [32] introduced prompt-based segmentation through point, box, text, or mask guidance. Recent studies [3, 33, 34, 35, 36] have demonstrated the strong adaptability of SAM across a variety of medical scenarios. These prompt-driven methods enabled flexible segmentation by incorporating domain knowledge directly into the inference process. To further refine segmentation outputs, PromptNucSeg [1] utilized an anchor-based detection prompter to automatically produce multi-point prompts within each nucleus for deeply guiding nuclei segmentation. UN-SAM [37] devised a multi-scale self-prompting strategy to extract coarse mask prompts from image embeddings, eliminating the demand for labor-intensive manual annotations. However, these prompt-based approaches are designed to improve performance within individual segmentation tasks. Unlike existing approaches that utilize prompts for task-specific enhancement, Co-Seg++ explores complementarities between semantic and instance segmentation in medical scenarios, which enables both tasks to mutually guide each other.
@@ 3 Methodology
We present the Co-Seg++ framework with the co-segmentation paradigm in Fig. 3, and achieve the mutual optimization between semantic and instance segmentation for medical images. To accomplish this, we design the STP-Encoder to provide prior spatial constraints to guide both tasks, and the MTC-Decoder to cooperatively generate semantic and instance maps through bidirectional information interactions, which enables the two tasks to enhance each other, leveraging spatio-tempral prompts for improved segmentation accuracy.
Figure 4: The illustration of STP-Encoder, integrating spatial and temporal prompts to establish long-range relationships between target segmentation regions and shared image embeddings, providing prior spatial constraints for cross-task guidance.
@@ 3.1 Co-Segmentation Paradigm
$\begin{split}p(y_{1},y_{2}|x,\theta_{1},\theta_{2})&=p(y_{1}|x,\theta_{1})\cdot p(y_{2}|y_{1},x,\theta_{2})\\ &=p(y_{2}|x,\theta_{2})\cdot p(y_{1}|y_{2},x,\theta_{1}).\end{split}$ (1)
Note that Eq. (1) reveals the mutual dependency between semantic and instance segmentation through intertwined conditional probabilities. It states that $p(y_{1},y_{2}|x,\theta_{1},\theta_{2})$, the joint probability of obtaining both segmentations given the image and model parameters, decomposed in two symmetrical ways:
  • $p(y_{1}|x,\theta_{1})\cdot p(y_{2}|y_{1},x,\theta_{2})$ suggests that once the semantic segmentation $y_{1}$ is known, it directly influences the prediction of the instance segmentation $y_{2}$.
  • Similarly, $p(y_{2}|x,\theta_{2})\cdot p(y_{1}|y_{2},x,\theta_{1})$ implies that knowing the instance segmentation $y_{2}$ affects the outcome of the semantic segmentation $y_{1}$.
@@
3.2 Spatio-Temporal Prompt Encoder
The existing prompt-based segmentation methods [32, 38] typically employ two distinct types of guidance signals: the sparse prompts (e.g., point and box) and the dense prompts (e.g., masks). While the sparse prompts provide precise region-of-interest cues, they offer limited contextual information for complex segmentation tasks compared to dense prompts. The dense prompts, in contrast, provide more comprehensive spatial cues but typically lack temporal context or task-specific structure. The proposed co-segmentation paradigm requires establishing implicit interdependencies between semantic and instance segmentation tasks. To achieve this, we introduce the STP-Encoder that leverages region masks and enriched image features to generate prior spatial constraints, effectively incorporating spatial and temporal prompts to improve the ability of the Co-Seg++ to coordinate between semantic and instance segmentation tasks.
As elaborated in Fig. 4, STP-Encoder comprises two parallel branches designed to capture temporal $t$ and task-specific spatial $s_{i}$ prompts. The temporal branch adopts shared image embeddings $h$ as input and first applies a linear layer $\rm Linear_{down}$ for down-projection. Then, a $3\times 3$ 1D convolution followed by a state space model (SSM) extracts temporal features and models long-range relationships. Finally, we apply another linear layer $\rm Linear_{up}$ for up-projection. These operations can be summarized as follows:
@@ 3.2 Spatio-Temporal Prompt Encoder
where $\rm LN(\cdot)$ denotes LayerNorm. Different from the existing Mamba networks [12, 24, 20], this temporal branch omits SiLU activation and gating operations since the image embeddings have already been optimally processed by the image encoder, avoiding redundant feature filtering. Moreover, the spatial branch employs several 2D convolutions to transform task-specific mask logits $m_{i}$ to patch embeddings, followed by self-attention $\rm SA(\cdot)$ to capture global context:
@@
[…]
@@
On this basis, we apply STP-Encoder to semantic and instance mask logits, effectively obtaining prior spatial constraints $c_{i}$ of Co-Seg++ for guiding versatile segmentation decoding.
@@
where $\oplus$ represents the element-wise addition. After that, we apply an MLP to refine the query embeddings on $q^{\prime}_{i}$, and further conduct the reverse cross-attention to generate task-specific image embeddings $g$, as follows:
@@
Dice mIoU HD Dice mIoU HD Dice mIoU HD Dice mIoU HD
@@
Dice mIoU HD Dice mIoU HD Dice mIoU HD
@@ 3.4 Optimization Pipeline
To construct our Co-Seg++ framework, we first adopt Hirea ViT [43] as the shared image encoder for versatile medical segmentation, ensuring the consistency of feature learning. In particular, we load [38] to initialize corresponding parameters and freeze these weights to preserve pre-trained knowledge. Additionally, we insert Adapter [44] into FFN layers to achieve parameter-efficient fine-tuning in medical domains. Based on the co-segmentation paradigm, Co-Seg++ includes two decoding forwards, where the semantic and instance heads of the MTC-Decoder share parameters in these two stages. For the first forward, MTC-Decoder leverages the image embedding generated from the image encoder to calculate binary segmentation masks: $y^{\rm bin}_{i}$ and $y^{\rm bin}_{j}$ of both tasks without any prompts. Then, we deliver them to the STP-Encoder to provide prior spatial constraints: $c_{i}$ and $c_{j}$ for both tasks. In the second forward, they will be directly transferred back to the MTC-Decoder to co-generate semantic and instance segmentation masks through cross-guidance. The training of Co-Seg++ consists of two parts: (1) the optimization of probability distribution and prior spatial constraints (2) the joint optimization of semantic and instance segmentation. The overall loss is formulated as:
@@
where $\lambda_{1}$ are factors to adjust the contribution of the first forward. By optimizing $\mathcal{L}_{\rm CoSeg}$, our framework enables accurate and consistent semantic and instance segmentation through joint learning with spatio-temporal priors and cross-task guidance. Co-Seg++ maintains computational efficiency by leveraging a shared encoder, reusing features across tasks, and introducing only lightweight adapter modules for medical adaptation. This collaborative design not only avoids redundant computation but also allows the two tasks to mutually reinforce each other, leading to more robust predictions. We now proceed to demonstrate the effectiveness of Co-Seg++ through comprehensive experiments across diverse datasets.

### v2

@@ Abstract
Medical image analysis is critical yet challenged by the need of jointly segmenting organs or tissues, and numerous instances for anatomical structures and tumor microenvironment analysis. Existing studies typically formulated different segmentation tasks in isolation, which overlooks the fundamental interdependencies between these tasks, leading to suboptimal segmentation performance and insufficient medical image understanding. To address this issue, we propose a Co-Seg++ framework for versatile medical segmentation. Specifically, we introduce a novel co-segmentation paradigm, allowing semantic and instance segmentation tasks to mutually enhance each other. We first devise a spatio-sequential prompt encoder (SSP-Encoder) to capture long-range spatial and sequential relationships between segmentation regions and image embeddings as prior spatial constraints. Moreover, we devise a multi-task collaborative decoder (MTC-Decoder) that leverages cross-guidance to strengthen the contextual consistency of both tasks, jointly computing semantic and instance segmentation masks. Extensive experiments on diverse CT and histopathology datasets demonstrate that the proposed Co-Seg++ outperforms state-of-the-arts in the semantic, instance, and panoptic segmentation of dental anatomical structures, histopathology tissues, and nuclei instances. The source code is available at https://github.com/xq141839/Co-Seg-Plus.
@@ 1 Introduction
Medical image segmentation plays a crucial role in clinical applications and has received significant research attention [1, 2, 3]. For example, histopathology image analysis demands accurate segmentation of tissue regions and separation of individual nuclei within these regions to assess tissue subtypes and tumor grading [4]. Similarly, cone beam computed tomography (CBCT) requires precise delineation of anatomical structures such as jaw bones and sinuses, alongside instance-level segmentation of individual teeth for dental treatment planning [5]. These dual requirements have given rise to the challenging field of versatile medical segmentation, which seeks to unify both semantic and instance segmentation tasks across various medical imaging contexts.
@@
Typically, medical image segmentation frameworks [6, 7, 8] have been designed for specific tasks, resulting in isolated solutions that address semantic segmentation and instance segmentation separately, as shown in Fig. 1(a). These conventional approaches typically adopt independent encoder-decoder architectures without shared information during the training process. The semantic segmentation networks are optimized to classify each pixel into pre-defined categories, meanwhile instance segmentation models [9, 10, 2] focus on detecting and delineating individual objects. However, multiple independent networks [11, 12, 13] lead to increased computation complexity. More seriously, this separation learning [14] impairs consistency for similar low-level features across tasks, resulting in redundant feature extraction. These limitations motivate the development of more integrated approaches that could leverage commonalities between semantic and instance segmentation tasks in medical images.
@@
To overcome this bottleneck, we propose Co-Seg++, a collaborative versatile medical segmentation framework that allows semantic and instance segmentation tasks to mutually enhance each other. As illustrated in Fig. 1(c), Co-Seg++, based on a novel co-segmentation paradigm, improves segmentation mask quality by capturing contextual dependencies between the two tasks. Specifically, we design a spatio-sequential prompt encoder (SSP-Encoder) that adaptively integrates spatial context and sequential dynamics to generate prior spatial constraints for guiding segmentation decoding across both tasks. We further introduce a multi-task collaborative decoder (MTC-Decoder) that leverages cross-task guidance to jointly predict semantic and instance segmentation masks in a unified framework. This collaborative learning strategy not only achieves enhanced contextual consistency but also significantly reduces prediction errors through mutual task optimization. Extensive experiments on diverse histopathology and dental datasets demonstrate that our Co-Seg++ consistently outperforms state-of-the-art segmentation methods.
@@
  • We devise a SSP-Encoder that integrates long-range spatial and sequential information to extract high-quality semantic and instance prompts, providing effective prior knowledge for cross-task segmentation guidance.
@@
Figure 3: The overview of our Co-Seg++ framework for collaborative versatile medical segmentation, consisting of SSP-Encoder that captures long-range spatial and sequential relationships and MTC-Decoder with semantic and instance heads for mutual task guidance. Co-Seg++ operates through two forward passes: the first pass generates binary masks from both heads as prior spatial constraints supervised by a spatial consistency constraint loss. In the second forward pass, the semantic and instance heads in MTC-Decoder leverage these constraints via cross-attention to mutually enhance semantic region delineation and instance classification. The Co-Seg++ framework fully exploits complementary contextual information through spatio-temporal prompts.
@@ 2.1 Medical Image Segmentation
Medical image segmentation is fundamental to computer-aided diagnosis and treatment planning [6, 7, 8]. This field encompasses semantic segmentation (i.e., pixel-level classification), instance segmentation (i.e., individual object delineation), and panoptic segmentation that unifies both by jointly predicting semantic classes and instance identities [19]. Building on the U-Net encoder-decoder framework [6], recent advances span three network architectures: convolutional neural network (CNN), vision transformers (ViT) [20], and more recent Mamba models [21]. The CNN-based methods [7, 22, 8] applied automated architecture optimization and multi-scale convolutions to enhance target awareness. The ViT-based methods [23, 24] leveraged self-attention mechanisms to capture global contexts. The Mamba-based methods [12, 25] utilized sequence modeling to handle long-range relationships with linear computational complexity. These works have demonstrated strong adaptability across diverse imaging modalities.
@@ 2.2 Versatile Segmentation Paradigms
The classical versatile medical segmentation frameworks have predominantly adopted a task-decoupled paradigm to address multiple segmentation objectives simultaneously. The TIAToolbox [14] employed an ensemble approach that integrates multiple independent networks for tissue and nuclei segmentation in histopathology images, treating each task as a separate optimization problem. To enhance cross-task consistency, Cerberus [15] employed the shared encoder with decoupled decoders for joint training efficiency. Federico et al. [30] adopted multi-class strategies with independent labels for CBCT tasks. VerSemi [31] introduced the task-prompted dynamic learning to provide the improved decoding workflow for different segmentation tasks. However, they lack flexible guidance mechanisms that limit their ability to incorporate contextual knowledge or interactive signals, highlighting the need for more adaptive and guidance-driven segmentation paradigms.
The segment anything model (SAM) [32] introduced prompt-based segmentation through point, box, text, or mask guidance. Recent studies [33, 34] have demonstrated the strong adaptability of SAM across a variety of medical scenarios. These prompt-driven methods enabled flexible segmentation by incorporating domain knowledge directly into the inference process. To further refine segmentation outputs, PromptNucSeg [1] utilized an anchor-based detection prompter to automatically produce multi-point prompts within each nucleus for deeply guiding nuclei segmentation. UN-SAM [35] devised a multi-scale self-prompting strategy to extract coarse mask prompts from image embeddings, eliminating the demand for labor-intensive manual annotations. However, these prompt-based approaches are designed to improve performance within individual segmentation tasks. Unlike existing approaches that utilize prompts for task-specific enhancement, Co-Seg++ explores complementarities between semantic and instance segmentation in medical scenarios, which enables both tasks to mutually guide each other.
@@ 3 Methodology
We present the Co-Seg++ framework with the co-segmentation paradigm in Fig. 3, and achieve the mutual optimization between semantic and instance segmentation for medical images. To accomplish this, we design the SSP-Encoder to provide prior spatial constraints to guide both tasks, and the MTC-Decoder to cooperatively generate semantic and instance maps through bidirectional information interactions, which enables semantic and instance segmentation tasks to enhance each other for improved segmentation accuracy.
Figure 4: The illustration of SSP-Encoder, integrating spatial prompts and sequential memory to establish long-range relationships between target segmentation regions and shared image embeddings, providing prior spatial constraints for cross-task guidance.
@@ 3.1 Co-Segmentation Paradigm
$\begin{split}p_{\theta_{1},\theta_{2}}(y_{1},y_{2}|x)&=p_{\theta_{1}}(y_{1}|x)\cdot p_{\theta_{2}}(y_{2}|y_{1},x)\\ &=p_{\theta_{2}}(y_{2}|x)\cdot p_{\theta_{1}}(y_{1}|y_{2},x).\end{split}$ (1)
Note that Eq. (1) reveals the mutual dependency between semantic and instance segmentation through intertwined conditional probabilities. It states that $p_{\theta_{1},\theta_{2}}(y_{1},y_{2}|x)$, the joint probability of obtaining both segmentations given the image and model parameters, is decomposed in two symmetrical ways:
  • $p_{\theta_{1}}(y_{1}|x)\cdot p_{\theta_{2}}(y_{2}|y_{1},x)$ suggests that once the semantic segmentation $y_{1}$ is known, it directly influences the prediction of the instance segmentation $y_{2}$.
  • Similarly, $p_{\theta_{2}}(y_{2}|x)\cdot p_{\theta_{1}}(y_{1}|y_{2},x)$ implies that knowing the instance segmentation $y_{2}$ affects the outcome of the semantic segmentation $y_{1}$.
@@
3.2 Spatio-Sequential Prompt Encoder
The existing prompt-based segmentation methods [32, 36] typically employ two distinct types of guidance signals: the sparse prompts (e.g., point and box) and the dense prompts (e.g., masks). While the sparse prompts provide precise region-of-interest cues, they offer limited contextual information for complex segmentation tasks compared to dense prompts. The dense prompts, in contrast, provide more comprehensive spatial cues but typically lack temporal context or task-specific structure. The proposed co-segmentation paradigm requires establishing implicit interdependencies between semantic and instance segmentation tasks. To achieve this, we introduce the SSP-Encoder that leverages region masks and enriched image features to generate prior spatial constraints, effectively incorporating spatial and sequential prompts to improve the ability of the Co-Seg++ to coordinate between semantic and instance segmentation tasks.
As elaborated in Fig. 4, SSP-Encoder comprises two parallel branches designed to capture temporal $t$ and task-specific spatial $s_{i}$ prompts. The temporal branch adopts shared image embeddings $h$ as input and first applies a linear layer $\rm Linear_{down}$ for down-projection. Then, a $3\times 3$ 1D convolution followed by a state space model (SSM) extracts sequential features and models long-range relationships. Finally, we apply another linear layer $\rm Linear_{up}$ for up-projection. These operations can be summarized as follows:
@@ 3.2 Spatio-Sequential Prompt Encoder
where $\rm LN(\cdot)$ denotes LayerNorm. Different from the existing Mamba networks [12, 24, 21], this sequential branch omits SiLU activation and gating operations since the image embeddings have already been optimally processed by the image encoder, avoiding redundant feature filtering. Moreover, the spatial branch employs several 2D convolutions to transform task-specific mask logits $m_{i}$ to patch embeddings, followed by self-attention $\rm SA(\cdot)$ to capture global context:
@@
where $\rm Conv2D_{1}$ denotes a $16\times 16$ convolution for patch tokenization, $\rm Conv2D_{2}$ is a $1\times 1$ convolution for channel expansion. After that, the SSP-Encoder adopts cross-attention to integrate spatial prompts with sequential prompts, as follows:
@@
On this basis, we apply SSP-Encoder to semantic and instance mask logits, effectively obtaining prior spatial constraints $c_{i}$ of Co-Seg++ for guiding versatile segmentation decoding.
@@ 3.3 Multi-Task Collaborative Decoder
where $\oplus$ represents the element-wise addition. When semantic features act as keys and values in the attention mechanism, while instance features serve as queries, this design allows the cross-attention operation to integrate global contextual information from semantic masks into instance-level features, enabling spatial constraints and shape priors to refine instance boundaries and reduce false positives. For example, closely packed or overlapping nuclei instances benefit from the broader spatial context provided by semantic features. After that, we apply a MLP to refine the query embeddings on $q^{\prime}_{i}$, and further conduct the reverse cross-attention to generate task-specific image embeddings $g$, as follows:
@@
[…]
@@
[…]
@@ 3.4 Optimization Pipeline
To construct our Co-Seg++ framework, we first adopt Hirea ViT [42] as the shared image encoder for versatile medical segmentation, ensuring the consistency of feature learning. In particular, we load [36] to initialize corresponding parameters and freeze these weights to preserve pre-trained knowledge. Additionally, we insert Adapter [43] into FFN layers to achieve parameter-efficient fine-tuning in medical domains. Based on the co-segmentation paradigm, the predictions for both tasks are based on shared image embeddings and mutual prompt guidance. Specifically, $\theta_{i}$ and $\theta_{j}$ contain the parameters of SSP-Encoder and the segmentation head of MTC-Decoder for different tasks. When task $j$ updates its parameters $\theta_{j}$, it affects the generation of spatial constraints $c_{j}$ in the first forward, which influence the prediction of task $i$ via the cross-attention mechanism in the second forward. This means that the gradient of task $i$ with respect to task $j$ parameters reflects how shared components allow one task to affect the other. This implicit gradient flow is critical for model performance, as it allows semantic segmentation and instance segmentation to mutually inform each other. In this way, semantic segmentation benefits from instance-level spatial constraints, while instance segmentation predictions are guided by global semantic context.
Moreover, the training of Co-Seg++ consists of two parts: (1) the optimization of probability distribution and prior spatial constraints (2) the joint optimization of semantic and instance segmentation. The overall loss is formulated as:
@@
where $\lambda_{1}$ are factors to adjust the contribution of the first forward. During backpropagation, the semantic loss in the second forward first influences the gradient updates of the image embeddings, semantic-specific query embeddings and instance-specific prior spatial constraints through the cross-attention mechanism. Specifically, the generation of instance-specific prior spatial constraints depends on binary instance masks that are generated by the image embedding and instance-specific query embeddings in the first forward. This creates a gradient flow pathway where the semantic loss propagates, through the instance-specific prior spatial constraints $c_{j}$, back to the instance query embeddings that generated the binary masks $y_{j}^{\rm bin}$ in the first forward. Therefore, the semantic loss directly influences the gradient updates of the instance-specific query embeddings, and vice versa, establishing the bidirectional optimization as described in our co-segmentation paradigm. By optimizing $\mathcal{L}_{\rm CoSeg}$, our framework enables accurate and consistent semantic and instance segmentation through joint learning with spatio-temporal priors and cross-task guidance. Co-Seg++ maintains computational efficiency by leveraging a shared encoder, reusing features across tasks, and introducing only lightweight adapter modules for medical adaptation. This collaborative design not only avoids redundant computation but also allows the two tasks to mutually reinforce each other, leading to more robust predictions. We now proceed to demonstrate the effectiveness of Co-Seg++ through comprehensive experiments across diverse datasets.